# Holographic duality from Howe duality: Chern–Simons gravity as an ensemble of code CFTs

### v1

@@ Appendix G Gap of the spectrum of $U(1)$ primaries
$\overline{W}=\frac{1}{2^{n/2}+4}\left(2^{n/2}(x^{n}+y^{n})+(x+y)^{n}+(x+iy)^{n}+(x-y)^{n}+(x-iy)^{n}\right)$ (241)
@@
$\overline{W_{\mathcal{C}}}\ =\ \frac{1}{1+p^{1-n}}\left(\Psi_{0}+p^{-n}\sum_{c\in{\mathscr{D}}}\Psi_{c}+p^{-n}\sum_{c\in{\mathscr{D}}}e^{\frac{\pi i\eta(c,c)}{p}}\Psi_{c}+\ldots+p^{-n}\sum_{c\in{\mathscr{D}}}e^{\frac{(p-1)\pi i\eta(c,c)}{p}}\Psi_{c}\right)$ (244)
@@
$N(n_{1},\ldots,n_{p-1})=\frac{(2n)!}{n_{0}!n_{1}!n_{2}!\ldots n_{p-1}!}$ (246)
@@
$\rho\sim\exp\left(2n\left(h_{p}-\frac{1}{2}\log p\right)\right)$ (248)
@@
$h_{p}$ $=\ -\log C+\mu p\delta$ (254)
$=\ \frac{1}{2}\log(2\pi p\delta)+\frac{1}{2}$ (255)

### v2

@@ Appendix G Gap of the spectrum of $U(1)$ primaries
$\overline{W}=\frac{1}{2^{n/2}+4}\left(2^{n/2}(x^{n}+y^{n})+(x+y)^{n}+(x+iy)^{n}+(x-y)^{n}+(x-iy)^{n}\right).$ (241)
@@
$\overline{W_{\mathcal{C}}}\ =\ \frac{1}{1+p^{1-n}}\left(\Psi_{0}+p^{-n}\sum_{c\in{\mathscr{D}}}\Psi_{c}+p^{-n}\sum_{c\in{\mathscr{D}}}e^{\frac{\pi i\eta(c,c)}{p}}\Psi_{c}+\ldots+p^{-n}\sum_{c\in{\mathscr{D}}}e^{\frac{(p-1)\pi i\eta(c,c)}{p}}\Psi_{c}\right).$ (244)
@@
$N(n_{1},\ldots,n_{p-1})=\frac{(2n)!}{n_{0}!n_{1}!n_{2}!\ldots n_{p-1}!}.$ (246)
@@
$\rho\sim\exp\left(2n\left(h_{p}-\frac{1}{2}\log p\right)\right),$ (248)
@@
$h_{p}\$$=\ -\log C+\mu p\delta$ (254)
$=\ \frac{1}{2}\log(2\pi p\delta)+\frac{1}{2}\,.$ (255)